OpenStudy (mikieandjj):

the radius of a cylinder is 3.5ft the height is 14 ft find the surface area and volume of the cylinder to the nearest tenth of a foot

OpenStudy (raffle_snaffle):

V_cylinder = pi*r^2*h pi = 3.14 r = radius h = height of cylinder

OpenStudy (raffle_snaffle):

Surface Area = 2(pi r 2) + (2 pi r)* h r = radius h = height of cylinder

OpenStudy (raffle_snaffle):

Okay tell me what you get?

OpenStudy (raffle_snaffle):



OpenStudy (raffle_snaffle):

radius is given to you and so is height just plug and chug. =)